# The TCF doesn’t really A(A)ID – Automatic Privacy Analysis and Legal Compliance of TCF-based Android Applications

### v1

@@ Abstract.
The Transparency and Consent Framework (TCF), developed by the Interactive Advertising Bureau (IAB) Europe, provides a de facto standard for requesting, recording, and managing user consent from European end-users. Its goal is to help organizations comply with the General Data Protection Regulation (GDPR) and the ePrivacy Directive (ePD). This framework has previously been found to infringe European data protection law and has subsequently been regularly updated. Previous research on the TCF focused exclusively on web contexts, with no attention given to its implementation in mobile applications. No work has systematically studied the privacy implications of the TCF on Android apps. To address this gap, we investigate the prevalence of the TCF in popular Android apps from the Google Play Store, and assess whether these apps respect users’ consent banner choices. By scraping and downloading 4482 of the most popular Google Play Store apps on an emulated Android device, we automatically determine which apps use the TCF, automatically interact with consent banners, and analyze the apps’ traffic in two different stages, passive (post choices) and active (during banner interaction and post choices).
We found that 576 (12.85%) of the 4482 downloadable apps in our dataset implemented the TCF, and we identified potential privacy violations within this subset. In 15 (2.6%) of these apps, users’ choices are stored only when consent is granted. Users who refuse consent are shown the consent banner again each time they launch the app. Network traffic analysis conducted during the passive stage reveals that 66.2% of the analyzed TCF-based apps share personal data, through the Android Advertising ID (AAID), in the absence of a lawful basis for processing. 55.3% of apps analyzed during the active stage share AAID before users interact with the apps’ consent banners, violating the prior consent requirement. These results suggest that the implementation of the TCF is prone to significant non-compliance practices, raising yet again serious concerns about its effectiveness in helping organizations adhere to regulatory requirements. We further expose concerns regarding Google as the dominant Consent Management Provider (CMP) in our dataset (89.76%), structurally accommodating potential legal violations.
Android, tracking, consent banner, privacy, compliance, TCF, legitimate interest, GDPR
@@ 1. Introduction
The majority of Web traffic now originates from mobile devices (similarweb), and in the past five years, Android has remained the most widely used mobile operating system (72.55% in October 2025) (mobile-os-marketshare). It has become increasingly common to rely on mobile apps – generally distributed through the Google Play Store –, to communicate and interact with devices and services, a phenomenon known as appification (case_appification_2019). However, the widespread use of these apps is accompanied by a large-scale tracking impacting users’ privacy (ap_news_google_2025).
@@
In the European Union (EU), personal data collection and processing is strictly regulated. The General Data Protection Regulation (GDPR) (GDPR2018) and the ePrivacy Directive (ePD) (ePD-09) specify data protection obligations for data controllers (herein mobile app publishers [^1] The French DPA (CNIL) qualifies app publishers as data controllers, and note that publishers and developers often merge in practice, although external developers can be considered a different contractual entity (cnil_recommendation_2025, p17,p7). and website owners) governing the lawful collection and processing of personal data. In particular, data protection law mandates that controllers identify and rely on a valid legal basis for each distinct purpose of processing, which may include consent or legitimate interest (LI) under Article 6 of the GDPR. However, regulatory practice and enforcement decisions have consistently affirmed that consent is the only legal basis for online behavioral advertising, instead of legitimate interest (CJEU2023_Bundeskartellamt; WP29_2010_OnlineBehaviouralAdvertising).
@@
To help organizations collect user consent according to the legal requirements, the Interactive Advertising Bureau (IAB) Europe – the European branch of an American advertising business organization –, introduced in April 2018 a policy and technical framework: the Transparency and Consent Framework (TCF) (TCF). The TCF is widely seen as the leading industry standard for managing online consent (waem_eu_2025), although the academic literature provides only anecdotal assessments of its adoption on the web (matte2020cookie; bollinger2022automating), and even less so on mobile contexts (koch2023ok). Although the TCF purports to help organizations meet data protection obligations, its deployment does not guarantee compliance. Prior analyzes have shown that websites relying on earlier TCF versions (i.e., v.1.1) nonetheless violated data protection law, despite formally implementing the framework (matte2020cookie; morel2023legitimate; Veale2022ImpossibleAsks; BelgianDPA2022_Decision21_2022). Specifically, the Belgian Data Protection Authority (APD) ruled in 2022 (BelgianDPA2022_Decision21_2022) that the TCF was non-compliant with the GDPR as it allowed the use of “legitimate interest” for targeted advertising – a legal basis deemed incompatible with the intrusive nature of large-scale behavioral tracking and profiling. The APD therefore required IAB Europe to remove “legitimate interest” as a valid ground for processing data for advertising purposes. Furthermore, the ruling concluded that the framework’s consent mechanism failed to meet the GDPR consent requirements, and subsequent rulings have been issued regarding the TCF (CJEU_IAB_TCF_2024).
@@
The framework has undergone several revisions since these rulings. An update (v.2.2) introduced in May 2023 required all TCF participants to implement the new version by November 2023. A major change introduced in this update was the removal of “legitimate interest” as a legal basis for processing personal data to create personalized advertising and content (TCF-v.2.2). While this reform was intended to align the framework more closely with GDPR compliance requirements, it remains unclear whether these obligations are consistently implemented and respected in practice.
@@ 1. RQ1
    How prevalent are TCF-based apps in the Google Play Store, and what are their characteristics?
@@ 2. RQ2
    Do TCF-based apps respect users’ consent choices and the TCF’s requirements of legitimate interest?
@@ 3. RQ3
To address these questions, we have set up a collaboration with a legal scholar in EU Data Protection law and we make the following contributions:
@@
    We measure the prevalence of the TCF in popular Android apps, including the first fine-grain quantification per category, and find a significant increase in its adoption, rising from 6.6%  reported in 2023 (koch2023ok) to 12.85% in our dataset.
@@
    We design a methodology to semi-automatically interact with consent banners deployed by major Consent Management Providers (CMPs) in TCF-based Android apps, combining three approaches capturing different choices to data collection (disagree to all, legitimate interest purposes only, agree to all).
@@
    We conduct the first systematic analysis of personal data processing in TCF-based Android apps through traffic analysis. Our results indicate that 15 apps stored consent choices only when provided with consent to all data processing, and that most apps (66.2%), with gaming apps as top violators, share personal data to ad servers against users’ choices. These results suggest a systemic breakdown in the TCF’s ability to facilitate legal compliance, calling into question its effectiveness as a privacy-governance mechanism.
@@
    We expose concerns regarding Google as the dominant CMP in our dataset (89.76%), structurally accommodating potential legal violations and offering bad defaults (pre-selected legitimate interests purposes).
@@ 2. Background knowledge
This section introduces background related to the GDPR and the ePD, AAID, and the TCF.
@@ 2.1. EU Data Protection Law
- P1 Lawfulness: when processing personal data, data controllers must collect personal data only with a valid legal basis (Art. 5(1)(a)), such as consent or legitimate interest (LI). Where the legal basis relied upon is consent – as the case when online tracking is carried out for advertising purposes, including in the context of this paper – the GDPR imposes strict requirements for valid consent. Consent must be i) prior to any data collection, ii) freely given, iii) specific, iv) informed, v) unambiguous, vi) readable, accessible, and vii) revocable (Articles 4(11) and 7 GDPR) (Sant-etal-20-TechReg). To be freely given, users must be able to accept or reject consent without facing negative consequences.
@@ 2.2. Google Advertising ID (AAID) as personal data
AAID as personal data under the GDPR. The online advertising industry, represented by IAB Europe, explicitly acknowledges that mobile advertising identifiers – including AAID – qualify as personal data under the GDPR (iab_europe_gdpr_implementation_working_group_definition_2018). This qualification is consistent with GDPR Recital 30, which states that identifiers provided by a user’s device, such as IP addresses or other device identifiers, may be used to single out individuals, thereby making them identifiable. The AAID mobile identifier is a direct example of such an “online identifier”. It is technically persistent, unique to the device, and is routinely used to track user behavior across apps and services. The French Data Protection Authority (CNIL) takes a similar stance, noting that mobile identifiers like the AAID fall under the scope of the ePrivacy Directive (cnil_recommendation_2025). This means their access and storage require user consent.
@@ 6.1.1. Growing prevalence of the TCF
We found 15 apps that only store user’ banner choices when the user gave consent to all data processing. When the user withheld consent for one or more purposes, the consent banner would reappear upon every subsequent launch of the app. Because the user’s choice is never registered, repeated prompts can pressure users to accept out of fatigue, potentially influencing their choices and bringing into question the freedom of consent expressed (cnil_recommendation_2025), which ultimately may infringe a freely given consent and thus the Lawfulness principle (P1) and the correct consent registration requirements (LR1). To this scope, such nagging practices, commonly identified as a dark pattern, appear closely tied to incorrect registration of consent banner choices (Sant-etal-20-TechReg).
@@ 6.3.1. Extending the dataset
Assessing the presence of the TCF on an app required downloading it, amounting to a significant overhead (87.15% of downloaded apps did not use the TCF). During interactions with consent banners, we noticed that many apps using the TCF were published by the same developers. In the Google Play Store, each application has a link to its developer, [^18] The Play Store refers to the developer, although the legal responsibility is first on the publisher (cnil_recommendation_2025). meaning that, given an app, a simple heuristic consisting in scraping all apps from the same developer can rapidly expand our dataset of TCF-based apps. We tried this method on a set of 842 applications initializing the framework, and found 4741 apps initializing the TCF too, all published by developers using the TCF in other apps. This method can bootstrap the creation of a larger dataset, enabling further privacy analysis, with the downside of not providing further data on proportions of TCF usage in the Play Store, and possibly biasing the representativeness of results. Another possibility of extension is to use a different sampling method, by downloading not only top apps but also long tail app with fewer downloads.
@@ 7. Conclusion
Acknowledgments
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and by the Utrecht Centre for Regulation and Enforcement in Europe (RENFORCE).

### v2

# arXiv:2602.20222v2[cs.CR] 27 Feb 2026

@@ Abstract.
The Transparency and Consent Framework (TCF), developed by the Interactive Advertising Bureau (IAB) Europe, provides a de facto standard for requesting, recording, and managing user consent from European end-users. Its goal is to help organizations comply with the General Data Protection Regulation (GDPR) and the ePrivacy Directive (ePD). This framework has previously been found to infringe European data protection law and has subsequently been regularly updated. Previous research on the TCF focused exclusively on web contexts, with no attention given to its implementation in mobile apps. No work has systematically studied the compliance implications of the TCF on Android apps. To address this gap, we investigate the prevalence of the TCF in popular Android apps from the Play Store, and assess whether these apps respect users’ consent banner choices. The TCF introduced minor changes in its new version (v.2.3) on March 1st, 2026, aimed at reducing ambiguity for vendors; these changes do not impact our results.
We scraped and downloaded 4482 of the most popular Google Play Store apps on an emulated Android device. We automatically identified that 576 (12.85%) of the 4482 downloadable apps implemented the TCF, and we detected potential legal violations within this subset. By automatically interacting with consent banners, we observed that in 15 (2.6%) of these apps, users’ choices are stored only when consent is granted. Users who refuse consent are shown the consent banner again each time they launch the app. We analyzed the apps’ traffic in two different stages, passive (post interaction with the banner) and active (during banner interaction and post user choices). Network analysis conducted during the passive stage reveals that 66.2% of the analyzed TCF-based apps share personal data through the Google Advertising ID (AAID) without consent – the lawful basis for such processing. Furthermore, 55.3% of apps analyzed during the active stage share AAID before users interact with the apps’ consent banners, violating the prior consent requirement. We further expose concerns regarding Google as the dominant Consent Management Provider (CMP) in our dataset (89.76%), structurally accommodating potential legal violations. Our results suggest that mobile implementations of the TCF are prone to significant non-compliance practices, raising concerns about its effectiveness in helping organizations adhere to legal requirements.
Android, online tracking, consent banner, online privacy, compliance, TCF v.2.2, legitimate interest, GDPR, ePrivacy Directive
@@ 1. Introduction
The majority of Web traffic now originates from mobile devices (similarweb), and in the past five years, Android has remained the most widely used mobile operating system (72.55% in October 2025) (mobile-os-marketshare). It has become increasingly common to rely on mobile apps – generally distributed through the Google Play Store –, to communicate and interact with devices and services, a phenomenon known as appification (case_appification_2019). However, the widespread use of these apps is accompanied by a large-scale tracking impacting users’ privacy (ap_news_google_2025). In the European Union (EU), personal data collection and processing is strictly regulated. The General Data Protection Regulation (GDPR) (GDPR2018) and the ePrivacy Directive (ePD) (ePD-09) specify legal requirements for data controllers (herein mobile app publishers [^1] The French DPA (CNIL) qualifies app publishers as data controllers, and note that publishers and developers often merge in practice, although external developers can be considered a different contractual entity (cnil_recommendation_2025, p17,p7). and website owners) to track end-users’s personal data. In particular, data protection law mandates that controllers identify a valid legal basis for each purpose of processing under Article 6 of the GDPR. Regulatory and enforcement decisions have consistently affirmed that consent is the only lawful basis for online behavioral advertising, instead of legitimate interest (CJEU2023_Bundeskartellamt; WP29_2010_OnlineBehaviouralAdvertising).
@@
To help organizations request user consent according to the legal requirements, the Interactive Advertising Bureau (IAB) Europe – the leading industry association for the digital marketing and advertising ecosystem in Europe – introduced in April 2018 the Transparency and Consent Framework (TCF) (TCF). The TCF is widely seen as the leading industry standard for managing online consent (waem_eu_2025), although academic literature demonstrates instances of its non-compliance practices on the Web  (matte2020cookie; bollinger2022automating; MatteSantosBielova2020; kancherla_johnny_2025), but even less so on mobile contexts (koch2023ok). Although the TCF purports to help organizations meet EU data protection law obligations, its deployment does not guarantee compliance. Prior analysis has shown that websites relying on earlier TCF versions (i.e., v.1.1) violated data protection law despite formally implementing the framework (matte2020cookie; morel2023legitimate; Veale2022ImpossibleAsks). Specifically, the Belgian Data Protection Authority (APD) ruled in 2022 (BelgianDPA2022_Decision21_2022) that the TCF was non-compliant with the GDPR as it allowed the use of “legitimate interest” for targeted advertising – a legal basis deemed incompatible with the intrusive nature of large-scale behavioral tracking and profiling. The APD therefore required IAB Europe to remove “legitimate interest” as a valid ground for processing data for advertising purposes. Furthermore, the ruling concluded that the framework’s consent mechanism failed to meet the GDPR consent requirements. The EU Court of Justice held that the TC String, used to record users’ consent preferences, can constitute personal data and it also found that IAB Europe may act as a joint controller with TCF participants even if it does not directly access the data (CJEU_IAB_TCF_2024).
@@
The framework has undergone several revisions since these rulings. The update (v.2.2) introduced in May 2023 required all TCF participants to implement the new version by November 2023. A major change introduced in this update was the removal of “legitimate interest” as a legal basis for processing personal data to create personalized ads and content (TCF-v.2.2). While this reform was intended to be legally compliant, it remains unclear whether the framework is consistently implemented and compliant. V.2.3 superseded v.2.2 on March 1st 2026, without incidence on our results.
@@ 1. RQ1
    How prevalent are TCF-based apps in the Google Play Store, and what are their characteristics (e.g., categories, country of origin, etc.)?
@@ 2. RQ2
    Do TCF-based apps respect users’ consent choices and the TCF’s requirements introduced by the v.2.2 not to use legitimate interest for certain purposes?
@@ 3. RQ3
To address these questions, we collaborated with a legal scholar in EU Data Protection law and we make the following contributions:
@@
    We measure the prevalence of TCF in popular Android apps, including the first fine-grain quantification per category, and find a significant increase in its adoption, rising from 6.6%  reported in 2023 (koch2023ok) to 12.85% in our dataset.
@@
    We design a methodology to semi-automatically interact with consent banners deployed by major Consent Management Providers (CMPs) in TCF-based Android apps, combining three approaches capturing different choices to data collection: “disagree to all” purposes, “legitimate interest purposes only”, and “consent to all”.
@@
    We perform the first systematic assessment of personal data processing in TCF-based Android apps through traffic analysis. Our results indicate that 15 apps stored consent choices only when provided with “consent to all” data processing, and that most apps (66.2%), with gaming apps as the top violators, share AAID to ad servers against users’ choices. We perform a legal analysis to assess compliance with GDPR and ePD requirements, which exposes several potential legal violations. These results suggest a systemic breakdown in the ability of the TCF to facilitate legal compliance, questioning its effectiveness as a privacy-governance mechanism.
@@
    We expose concerns regarding Google as the dominant CMP in our dataset (89.76%), structurally accommodating potential legal violations, and offering bad defaults (pre-selected legitimate interest-based purposes).
@@ 2. Background knowledge
This section introduces background related to GDPR and the ePD, AAID, and TCF.
@@ 2.1. EU Data Protection Law
- P1 Lawfulness: when processing personal data, data controllers must collect personal data only with a valid legal basis (Art. 5(1)(a)), such as consent or legitimate interest (LI). In the context of online tracking for (non-essential) advertising purposes, consent is the lawful legal basis according to Article 5(3) of the ePrivacy Directive. Further, the GDPR imposes strict requirements for valid consent. Consent must be i) prior to any data collection, ii) freely given, iii) specific, iv) informed, v) unambiguous, vi) readable, accessible, and vii) revocable (Articles 4(11) and 7 GDPR) (Sant-etal-20-TechReg). To be freely given, users must be able to accept or reject consent without facing negative consequences.
@@ 2.2. Google Advertising ID (AAID) as personal data
AAID as personal data. The online advertising industry, represented by IAB Europe, explicitly acknowledges that mobile advertising identifiers – including AAID – qualify as personal data under the GDPR (iab_europe_gdpr_implementation_working_group_definition_2018). This qualification is consistent with GDPR Recital 30, which states that identifiers provided by a user’s device, such as IP addresses or other device identifiers, may be used to single out individuals, thereby making them identifiable. The AAID mobile identifier is a direct example of such an “online identifier”. It is technically persistent, unique to the device, and is routinely used to track user behavior across apps and services. The French Data Protection Authority (CNIL) takes a similar stance, noting that mobile identifiers like the AAID fall under the scope of Article 5(3) of the ePrivacy Directive (cnil_recommendation_2025). This means their access and storage require user consent.
@@ 6.1.1. Growing prevalence of the TCF
We found 15 apps that only store user’ banner choices when the user gave consent to all data processing. When the user withheld consent for one or more purposes, the consent banner would reappear upon every subsequent launch of the app. Because the user’s choice is never registered (as demanded by the accountability principle under Article 7(1) of the GDPR), repeated prompts can pressure users to accept out of fatigue, potentially influencing their choices and bringing into question the freedom of consent expressed (cnil_recommendation_2025), which ultimately may infringe a freely given consent and thus the Lawfulness principle (P1), the correct consent registration requirements (LR1), and the principle of Data Protection by Design (P5). To this scope, such nagging practices, commonly identified as a dark pattern, appear closely tied to incorrect registration of consent banner choices (Sant-etal-20-TechReg).
@@ 6.3.1. Extending the dataset
Assessing the presence of the TCF on an app required downloading it, amounting to a significant overhead (87.15% of downloaded apps did not use the TCF). During interactions with consent banners, we noticed that many apps using the TCF were published by the same developers. In the Google Play Store, each application has a link to its developer, [^18] The Play Store refers to the developer, although the legal responsibility is first on the publisher (cnil_recommendation_2025). meaning that, given an app, a simple heuristic consisting in scraping all apps from the same developer can rapidly expand our dataset of TCF-based apps. We tried this method on a set of 842 applications initializing the framework, and found 4741 apps initializing the TCF too, all published by developers using the TCF in other apps. This method can bootstrap the creation of a larger dataset, enabling further privacy analysis, with the downside of not providing further data on proportions of TCF usage in the Play Store, and possibly biasing the representativeness of results. Another possibility of extension is to use a different sampling method, by downloading not only top apps but also long tail apps with fewer downloads.
@@ 7. Conclusion
Acknowledgements.
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and by the Utrecht Centre for Regulation and Enforcement in Europe (RENFORCE). The authors would like to thank the reviewers for helping us improve this paper.